OpenStudy (anonymous):

Help pleaseeee Which of the following ordered pairs is a solution to the inequality y -2x + 10? (0, -4) (1, 13) (6, -1) (3, 8)

jimthompson5910 (jim_thompson5910):

the inequality sign is missing

jimthompson5910 (jim_thompson5910):

what's between the y and -2x ?

OpenStudy (anonymous):

oops it is <=

jimthompson5910 (jim_thompson5910):

what you can do is check each possible answer by plugging it into the inequality

OpenStudy (anonymous):

to see if its true or false?

jimthompson5910 (jim_thompson5910):

so let's check choice D (just a random answer choice I picked) in choice D, we have (3,8) which means x = 3 and y = 8 y <= -2x + 10 8 <= -2(3) + 10 ... plug in x = 3 and y = 8, now simplify 8 <= -6 + 10 8 <= 4 ... FALSE Since 8 <= 4 is FALSE, this means y <= -2x + 10 is FALSE when x = 3 and y = 8 So (3,8) is NOT a solution So choice D is eliminated as a possible answer

jimthompson5910 (jim_thompson5910):

yes exactly, you're looking for points that make y <= -2x + 10 true

OpenStudy (anonymous):

ohh okay i think i get it now @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking, tell me what you get

OpenStudy (anonymous):

for A i got -4<=18

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

good, that is correct and it is a true statement so choice A is a solution

jimthompson5910 (jim_thompson5910):

luckily this sounds like it's a question where only one answer is allowed if that's true, then you can stop once you found A is the answer

jimthompson5910 (jim_thompson5910):

but sometimes there will be other answers as well, so it doesn't hurt to check them all

OpenStudy (anonymous):

Thank you for the help (: @jim_thompson5910 and i did B and C too

jimthompson5910 (jim_thompson5910):

ok great, you'll find that B and C don't work, so they're eliminated you're welcome